in a series circuit, the voltage drop across a resistor is directly proportional to the resistance. if the voltage drops across a 75 ohm resistance is 10 volts, what is the voltage drop across a 30 ohm resistor?
Answer by stanbon(75887)

in a series circuit, the voltage drop across a resistor is directly proportional to the resistance. if the voltage drops across a 75 ohm resistance is 10 volts, what is the voltage drop across a 30 ohm resistor?
-----
v = k*r
---
Solve for "k" using "the voltage drops across a 75 ohm resistance is 10 volts"
10 = k*75
k = 2/15
------
Equation:
v = (2/15)*r
----
What is the voltage drop across a 30 ohm resistor?
v = (2/15)*30
v = 4 volts drop
===================
Cheers,
Stan H.